Determine the electric potential at the center of a square if 4C charges
are placed at the vertices of the square of side length 5mm
1
Expert's answer
2021-08-31T17:04:59-0400

Given

Charge of each charge is Q= 4 C

Length of square is "L=5 \times 10^{-3} m"

The electric potential at the centre of the square is

"V=\frac{4kQ}{\frac{L}{\sqrt 2}}=\sqrt 2\frac{4kQ}{L}"

"=\sqrt 2\frac{4(9\times 10^9)(4 )}{5 \times 10^{-3}}=4.07\times 10^{13} V"